# Weak instrumental variables due to nonlinearities in panel data: A Super Learner Control Function estimator

### v3

@@ 11.3 Assumptions
The set $\mathcal{G}$ consists P-integrable functions $\mathbf{T}g$ with pth moment that exists and it contains $\mathbf{T}g_{o}$, and
$\left\|\mathbf{T}g-\mathbf{T}g_{o}\right\|_{P,2}\leq\delta_{N}.$
@@
    $\mathbb{E}[\mathbb{E}_{n_{T,b}}\phi(O_{i};\theta_{o},\mathbf{T}g_{o})]=0.$
@@
    $\mathbb{E}[\mathbb{E}_{n_{T,b}}\phi(O_{i};\theta_{o},\mathbf{T}g_{o})\phi(O_{i};\theta_{o},\mathbf{T}g_{o})^{\prime}]=\Sigma_{i}$ with $\Sigma_{i}$ a semi positive-definite matrix, with eigenvalues bounded away from 0 and finite, with the minimum eigenvalue $\lambda_{T}>0$, and the maximum eigenvalue $\lambda_{1}<\infty$.
@@
    There exists finite real constants $C_{1}$,$C_{2}$, $C_{3}$, $C_{4}$ satisfying $||y_{i}||_{P,p}\leq C_{1}$, $||z_{i}||_{P,p}\leq C_{2}$, $||x_{1,i}||_{P,p}\leq C_{3}$, $||\tilde{x}_{i}||_{P,p}\leq C_{4}$ for all $k>2$.
@@ 11.4 Consistency
$\begin{split}\phi(O_{i};\theta,\mathbf{T}g)-\phi(O_{i};\theta,\mathbf{T}g_{o})&=(\mathbb{T}_{i}y_{i}-\tilde{H}_{i}\theta)^{\prime}V_{i}^{-1}\tilde{H}_{i}-(\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}H_{i}\theta)^{\prime}V_{i}^{-1}\mathbb{T}_{i}H_{i}\\ &=(\mathbb{T}_{i}y_{i})^{\prime}V_{i}^{-1}[\tilde{H}_{i}-\mathbb{T}_{i}H_{i}]+[(\tilde{H}_{i}-\mathbb{T}_{i}H_{i})\theta)]^{\prime}V_{i}^{-1}[(\tilde{H}_{i}-\mathbb{T}_{i}H_{i})],\end{split}$ (38)
Now, notice that $(\tilde{H}_{i}-\mathbb{T}_{i}H_{i})=[\mathbb{T}_{i}x_{1i},\mathbb{T}_{i}\tilde{X}_{i},\widehat{\mathbf{T}u}_{i}]-[\mathbb{T}x_{1i},\mathbb{T}\tilde{X}_{i},\mathbb{T}u_{i}]=[0,0,\widehat{\mathbf{T}u}_{i}-\mathbb{T}_{i}u_{i}]=[0,0,\xi_{i}]$. Then, we are left with:
$(\mathbb{T}_{i}y_{i})^{\prime}V_{i}^{-1}\xi_{i}+\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}.$ (39)
@@
$\begin{split}\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\mathbb{E}_{n_{T,b}}[(\mathbb{T}_{i}y_{i})^{\prime}V_{i}^{-1}\xi_{i}+\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}]]\right\|&\leq\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[(\mathbb{T}_{i}y_{i})^{\prime}V_{i}^{-1}\xi_{i}]\right\|+\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\sum_{t}\lambda_{t}(q_{t}^{\prime}\mathbb{T}_{i}y_{i})(q_{t}^{\prime}\xi_{i})]\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\theta\mathbb{E}_{P}[\sum_{t}\lambda_{t}(q_{t}^{\prime}\xi_{i}^{\prime})(q_{t}^{\prime}\xi_{i})]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\lambda_{1}\sum_{t}(q_{t}^{\prime}\mathbb{T}_{i}y_{i})(q_{t}^{\prime}\xi_{i})]\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}[\lambda_{1}\left\|\theta\mathbb{E}_{P}\sum_{t}(q_{t}^{\prime}\xi_{i})(q_{t}^{\prime}\xi_{i})]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\left\|\lambda_{1}\left\|\mathbb{T}_{i}y_{i}\right\|_{P,2}\left\|\xi_{i}\right\|_{P,2}\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\mathbf{T}g\in\mathcal{G}}\lambda_{1}\left\|\theta\right\|\left\|\xi_{i}\right\|_{P,2}\end{split}$ (40)
Since $\Theta$ is bounded, $\left\|\mathbf{T}g-\mathbf{T}g_{o}\right\|_{P,2}<\delta_{N}$, the last expression is bounded by $\delta_{N}$ by assumptions 5, and 7.

### v4

@@ 11.3 Assumptions
The set $\mathcal{G}$ consists P-integrable functions $\tau g$ with pth moment that exists and it contains $\tau g_{o}$, and
$\left\|\tau g-\tau g_{o}\right\|_{P,2}\leq\delta_{N}.$
@@
    $\mathbb{E}[\mathbb{E}_{n_{T,b}}\phi(O_{i};\theta_{o},\tau g_{o})]=0.$
@@
    $\mathbb{E}[\mathbb{E}_{n_{T,b}}\phi(O_{i};\theta_{o},\tau g_{o})\phi(O_{i};\theta_{o},\tau g_{o})^{\prime}]=\Sigma_{i}$ with $\Sigma_{i}$ a semi positive-definite matrix, with eigenvalues bounded away from 0 and finite, with the minimum eigenvalue $\lambda_{T}>0$, and the maximum eigenvalue $\lambda_{1}<\infty$.
@@
    There exists finite real constants $C_{1}$$C_{2}$, $C_{3}$, $C_{4}$ satisfying $y_{i}||_{P,p}\leq C_{1}$, $z_{i}||_{P,p}\leq C_{2}$, $x_{1,i}||_{P,p}\leq C_{3}$, $\tilde{x}_{i}||_{P,p}\leq C_{4}$ for all $k>2$.
@@ 11.4 Consistency
$\begin{split}\phi(O_{i};\theta,\tau g)-\phi(O_{i};\theta,\tau g_{o})&=(M_{\tau_{i}}y_{i}-\tilde{H}_{i}\theta)^{\prime}V_{i}^{-1}\tilde{H}_{i}-(M_{\tau_{i}}y_{i}-M_{\tau_{i}}H_{i}\theta)^{\prime}V_{i}^{-1}M_{\tau_{i}}H_{i}\\ &=(M_{\tau_{i}}y_{i})^{\prime}V_{i}^{-1}[\tilde{H}_{i}-M_{\tau_{i}}H_{i}]+[(\tilde{H}_{i}-M_{\tau_{i}}H_{i})\theta)]^{\prime}V_{i}^{-1}[(\tilde{H}_{i}-M_{\tau_{i}}H_{i})],\end{split}$ (38)
Now, notice that $(\tilde{H}_{i}-M_{\tau_{i}}H_{i})=[M_{\tau_{i}}x_{1i},M_{\tau_{i}}\tilde{X}_{i},\widehat{\tau u}_{i}]-[M_{\tau_{i}}x_{1i},M_{\tau_{i}}\tilde{X}_{i},M_{\tau_{i}}u_{i}]=[0,0,\widehat{\tau u}_{i}-M_{\tau_{i}}u_{i}]=[0,0,\xi_{i}]$. Then, we are left with:
$(M_{\tau_{i}}y_{i})^{\prime}V_{i}^{-1}\xi_{i}+\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}.$ (39)
@@
$\begin{split}\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\mathbb{E}_{n_{T,b}}[(M_{\tau_{i}}y_{i})^{\prime}V_{i}^{-1}\xi_{i}+\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}]]\right\|&\leq\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[(M_{\tau_{i}}y_{i})^{\prime}V_{i}^{-1}\xi_{i}]\right\|+\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\theta^{\prime}\xi_{i}^{\prime}V_{i}^{-1}\xi_{i}]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\sum_{t}\lambda_{t}(q_{t}^{\prime}M_{\tau_{i}}y_{i})(q_{t}^{\prime}\xi_{i})]\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\theta\mathbb{E}_{P}[\sum_{t}\lambda_{t}(q_{t}^{\prime}\xi_{i}^{\prime})(q_{t}^{\prime}\xi_{i})]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\mathbb{E}_{P}[\lambda_{1}\sum_{t}(q_{t}^{\prime}M_{\tau_{i}}y_{i})(q_{t}^{\prime}\xi_{i})]\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}[\lambda_{1}\left\|\theta\mathbb{E}_{P}\sum_{t}(q_{t}^{\prime}\xi_{i})(q_{t}^{\prime}\xi_{i})]\right\|\\ &\leq\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\left\|\lambda_{1}\left\|M_{\tau_{i}}y_{i}\right\|_{P,2}\left\|\xi_{i}\right\|_{P,2}\right\|+\\ &\operatorname*{sup}_{\theta\in\Theta,\tau g\in\mathcal{G}}\lambda_{1}\left\|\theta\right\|\left\|\xi_{i}\right\|_{P,2}\end{split}$ (40)
Since $\Theta$ is bounded, $\left\|\tau g-\tau g_{o}\right\|_{P,2}<\delta_{N}$, the last expression is bounded by $\delta_{N}$ by assumptions 5, and 7.